I am taking a biochemistry class and it has been a long time since I've had to work with logs/Logs (I don't understand the difference). This is my question:
160x10^-9=Log160^-9=2.2^-9=-6.8
They had me up to 2.2^-9. How does that become -6.8?

Answer by scott8148(6628)

multiplying quantities can be done by adding their logs

log(a * b) = log(a) + log(b)

log(160x10^-9) = log(160) + log(10^-9) = 2.2 + (-9) = -6.8